Operations Report – September 13th

MDRS Complete Operations Report 13-SEPT-2021

Name of person filing report: Shannon Rupert

Reason for Report: Routine daily

Non-nominal systems: Nothing to report

Action taken for non-nominal systems: Nothing to report

Generator: Nothing to report

ScienceDome Dual Split: Operating

Solar— Nominal, providing all power

Solar— SOC % Last 24 hours:

Average 75.3

Minimum 52

Maximum 100

Note on solar: We did not use the upper deck fan today. Apparently it uses a lot of power.

Diesel Reading – Empty, tank to be removed soon

Propane Reading, station tank – 80%

Propane Reading, director tank— 75%

Propane Reading, intern tank— 72%

Propane Reading, generator: 80%

Ethanol Free Gasoline – 5 gallons (and it looks like we are missing a couple of gas cans)

Water (loft tank): 50 gallons

Water Meter: n/a units

Water (static tank) – 250 gallons

Static to Loft Pump used – yes

Water in GreenHab – 70 gallons

Water in ScienceDome: 0 gallons

Water (Outpost tank) – 250 gallons

Hab toilet tank emptied: no

Perseverance rover used: Yes

Hours: 219.4

Beginning charge: 100

Ending charge: 100

Currently charging: yes

Sojourner rover used: no

Hours:

Beginning charge: 100

Ending charge: 100

Currently charging: yes

Spirit rover used: no

Hours: 136.9

Beginning charge: 100

Ending charge: 100

Currently charging: yes

Opportunity rover used: no

Hours: 72.3

Beginning charge: 100

Ending charge: 95

Currently charging: no

Curiosity rover used: no

Hours: 144.4

Beginning charge: 100

Ending charge: 100

Currently charging: yes

Notes on rovers: Nothing to report

ATV’s Used: (Honda, 300, 350.1, 350.2, 350.3): Honda

Reason for use: put new battery in, currently running fine but needs air in rear tires

Oil Added? No

ATV Fuel Used: 5 Gals

# Hours the ATVs were Used today: 0

Notes on ATVs: 350.2 would not start today, low battery

HabCar used and why, where? Yes, garbage and water

CrewCar used and why, where? No

Luna used and why, where? no

General notes and comments: Nothing to report

Summary of internet: All nominal.

Summary of suits and radios: Nothing to report

Campus wide inspection, if action taken, what and why? Low area outside SciDome that always has a puddle has been filled with sand

Summary of general operations: Nothing to report

Summary of Hab operations: Lounge chairs removed from upper deck, replaced by small sofa until new lounge couch is back in stock. Medicine cabinet from shower room was relocated to the main room to serve as our first aid kit for the Hab.

Summary of Outpost operations: Nothing to report

Summary of GreenHab operations: Filled the remaining carrot bin with soil.

Summary of ScienceDome operations: Nothing to report

Summary of RAM operations: Nothing to report

Summary of any observatory issues: Nothing to report

Summary of health and safety issues: Nothing to report

Questions, concerns, supplies needed and requests: Any updates on possible care products for the rover’s plastic parts?

Operations Report – September 12th

MDRS Complete Operations Report 12-SEPT-2021

Name of person filing report: Shannon Rupert

Reason for Report: Beginning of field season/work party daily update

Non-nominal systems: Nothing to report

Action taken for non-nominal systems: Nothing to report

Generator: Nothing to report

ScienceDome Dual Split: Operating

Solar— Nominal, providing all power

Solar— SOC % Last 24 hours:
Average 68.3
Minimum 30
Maximum 100
Note on solar: Nothing to report

Diesel Reading – 0 %.

Propane Reading, station tank – 80 %
Propane Reading, director tank— 75%
Propane Reading, intern tank— 72%
Propane Reading, generator: 80%

Ethanol Free Gasoline – 5 gallons (and it looks like we are missing a couple of gas cans)

Water (loft tank): 50 gallons

Water Meter: n/a units

Water (static tank) – under 100 gallons

Static to Loft Pump used – yes

Water in GreenHab – 90 gallons

Water in ScienceDome: 0 gallons

Water (Outpost tank) – 150 gallons

Hab toilet tank emptied: no

Perseverance rover used: Yes
Hours: 219.4
Beginning charge: 100
Ending charge: 100
Currently charging: yes

Sojourner rover used: yes
Hours:
Beginning charge: 100
Ending charge: 100
Currently charging: yes

Spirit rover used: yes
Hours: 136.9
Beginning charge: 100
Ending charge: 100
Currently charging: yes

Opportunity rover used: no
Hours: 72.3
Beginning charge: 100
Ending charge: 95
Currently charging: no

Curiosity rover used: yes
Hours: 144.4
Beginning charge: 100
Ending charge: 100
Currently charging: yes

Notes on rovers: Nothing to report

ATV’s Used: (Honda, 300, 350.1, 350.2, 350.3): 350.3
Reason for use: recreation
Oil Added? No
ATV Fuel Used: 5 Gals

# Hours the ATVs were Used today: 0.5
Notes on ATVs: Nothing to report

HabCar used and why, where? Yes, unhitched power washer and loaded to take garbage to town today

CrewCar used and why, where? No

Luna used and why, where? Yes, to Grand Junction for supplies

General notes and comments: Holes were dug for the new signs, however, we were unable to purchase the pier blocks needed for the completion of this project.

Summary of internet: All nominal.

Summary of suits and radios: Nothing to report

Campus-wide inspection, if action is taken, what and why? Nothing to report

Summary of general operations: Nothing to report

Summary of Hab operations: Water tank and cleaned and repainted with a thick layer of elastomeric paint. The bottom had deteriorated and will need to be repaired. I have a plan to replace the lower inch or so with foam tubbing and then repaint.

Summary of Outpost operations: Water tank and cleaned and repainted with a thick layer of elastomeric paint. This tank in particular had a minimum coat of paint with many spots where the foam was exposed, so we are hoping this will protect it. The tank also needs the bottom inch or so redone.

Summary of GreenHab operations: The floor was power washed.

Summary of ScienceDome operations: Nothing to report. If the weather holds, later this week we will do a final coat on the lower portion to create a thicker coat of paint there.

Summary of RAM operations: Nothing to report

Summary of any observatory issues: Nothing to report

Summary of health and safety issues: The cement pad in the back of the Hab cracked a long time ago but this summer the rains lifted in half and there is about a one-inch lift on one side. Not much can be done but warning paint will be applied.

Questions, concerns, supplies needed and requests: Nothing at this time.